What uniform magnetic field must be set up in space to permit a proton of speed 1.0x10^7 m/s to move in a circle the size of the earth’s equator? (R = 6400 km) 


1
Expert's answer
2021-11-03T11:18:56-0400

We know that

"R=\frac{mv}{qB}"

"B=\frac{mv}{qR}"


"B=\frac{1.6\times10^{-27}\times10^7}{1.6\times10^{-19}\times64\times10^5}=1.63\times10^{-8}T"